0 5 18 43 84 145 ? what is the next number in this given series?

Answer by Edwin McCravy(20079)

Write the numbers in a column:

  0
  5 
 18 
 43 
 84 
145

Starting with the first number 0, subtract it from
the number just below it, 5, getting 5 and write the
difference out to the right of that number in a 
second column. Do that all the way down to 185-84=61

  0   5     
  5  13         
 18  25     
 43  41       
 84  61    
145        

Do the same thing with the second column. Starting with 
the first number 5, subtract it from the number just 
below it, 13, getting 8, and write the difference
out to the right of that number in a third column.
Do that all the way doen to 61-41=20

  0   5   8     
  5  13  12         
 18  25  16      
 43  41  20     
 84  61    
145

Make one more column in the same way, and we see
that they are all 4's

  0   5   8   4  
  5  13  12   4       
 18  25  16   4   
 43  41  20     
 84  61    
145

Since they are all 4's we put another 4 at the 
bottom of the fourth column and work backward 
from right to left there by adding.  That is 
4+20=24, 24+61=85, 85+145=230

  0   5   8   4  
  5  13  12   4       
 18  25  16   4   
 43  41  20   4     
 84  61  24    
145  85
230

Answer: 230.

Edwin